OpenStudy (anonymous):

write the equation of the line that is perpendicular to y= 1/4x and passes through the point (2,-5)

OpenStudy (anonymous):

ok .. how do i find the equation

OpenStudy (anonymous):

(y-5)=4(x--2) --->y-5=4x+8 --->y=4x+13

OpenStudy (anonymous):

is this the answer

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

y=4x+13 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Any doubts??:)

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

Your answer is wrong just tried it.

OpenStudy (anonymous):

what is the answer then?

OpenStudy (anonymous):

it doesnt tell you it. Unless you get it right

OpenStudy (anonymous):

let me ask you a ques..is this y=1/(4x)

OpenStudy (anonymous):

opps i Made a mistake in the problem Its supposed to be (2,-5) My bad I'm just so tired right now

OpenStudy (anonymous):

and no its like this .. y = 1/4x

OpenStudy (anonymous):

ok..so then it is y--5=4(x-2) =y+5=4x-8 -->y=4x-3

OpenStudy (anonymous):

try this

OpenStudy (anonymous):

still wrong /:

OpenStudy (anonymous):

ya now i made a mistake

OpenStudy (anonymous):

it is y--5=-4(x-2) --->y+5=-4x+8 --->y=-4x+3

OpenStudy (anonymous):

is this the right one?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

still wrong..

OpenStudy (anonymous):

oh come on..

OpenStudy (anonymous):

whats happening..

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

the problem is right.

OpenStudy (anonymous):

y=(1/4)x are you sure and is it 'perpendicular'?

OpenStudy (anonymous):

not parallel

OpenStudy (anonymous):

IM SURE its perpendicular

OpenStudy (anonymous):

i read over it plenty of times

OpenStudy (anonymous):

do you know how to do this ..

OpenStudy (anonymous):

this is how it is done..i think it should be given in standard form

OpenStudy (anonymous):

4x+y-3=0

OpenStudy (anonymous):

same thing..different appearence

OpenStudy (anonymous):

will you try with this

OpenStudy (anonymous):

@xoxox123

OpenStudy (anonymous):

its wrong maybe you made a mistake in your math ?

OpenStudy (anonymous):

lol I've tried it like 10 times already -_-

OpenStudy (anonymous):

Dang...i cant find whats wrong here :(

OpenStudy (anonymous):

ohhh. well its okay.. then. thanks anyway

OpenStudy (anonymous):

for trying